A researcher wants to know if the paint color of a room influences the emotional experience of anger. To examine this, he splits people into two equal group and asks them to write about an experience that made them angry. They then rated on a scale from 1-10 how angry they were at that moment. Half of the participants did this in a room painted blue and the other half did this in a room painted red. The researcher reported a t (48) = 1.56, p > .05 for an independent-measures experiment. ​ How many individuals participated in the condition or group where they recalled an angry memory in the red room?
Answer #1

Degrees of freedom = n1 + n2 - 2 = 48

Hence,

n1 + n2 = 48 + 2 = 50

So,

Number of individuals participated in the condition or group where they recalled an angry memory in the red room = 50